Grade 12th passMechanics

Calculate the hydrostatic difference in blood pressure in a person of height 1.83 m between the brain and foot

Nisar Ahmed Dars
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Nisar
 
you have to use Stevin's law 
ΔP=d*g*Δh 
ΔP = 1060 kg/m^3 *9.8 m/s^2 *1.83 m= 19,010.04 Pa (kg*m*s^-2/m^2 = N/m^2 = Pa)
 
 
Regards
Arun (askIITians forum expert)